# Classification of three-qubit genuine entangled states using concurrence fill

### v1

@@ IV Concurrence Fill for a Rank-two Mixture of states
In this section, we study the rank-2 mixture of generalized GHZ and W states. Firstly we derive the minimum value of the concurrence fill for any pure state superposition of generalized GHZ and W states. Next, we formulate the concurrence fill for the mixed states using optimal decomposition.
@@ IV.1 Concurrence fill for eigenstates of $\rho(p)$
In order to find the concurrence fill for mixed states, let us first study the pure states that are superpositions of the eigenstates of $\rho$. Note that these states belong to the basis given in (15).
$|\psi_{p,\phi}\rangle=\sqrt{p}|gGHZ\rangle-\sqrt{1-p}\;e^{i\phi}|gW\rangle$ (35)
Three-tangle for the pure state $|\psi_{p,\phi}\rangle$ is given as follows oster2006 ; oster2008 ; njp2008 .
@@
Using (12), we derive the concurrence fill of the state $|\psi_{p,\phi}\rangle$ and it is given as follows:
@@
$C_{F}(|gGHZ\rangle)$ $=$ $4|a^{2}b^{2}|$ (39)
$C_{F}(|gW\rangle)$ $=$ $8|cdf|\left(\frac{|c^{2}d^{2}+d^{2}f^{2}+c^{2}f^{2}|}{3}\right)^{1/4}$ (40)
Now we derive a lower bound of the concurrence fill for the pure eigenstates of the form (35), which is appended in the following result.
Result. For the three-qubit pure state $|\psi_{p,\phi}\rangle$ given in (35), the following inequality holds.
$C_{F}(|\psi_{p,\phi}\rangle)\geq C_{F}^{(min)}(|\psi_{p_{min},\phi}\rangle)$ (41)
where $p_{min}=\frac{1}{1+3b^{2}}$, and
@@
represents the lower bound of the concurrence fill for the state $|\psi_{p,\phi}\rangle$.
@@
In order to derive the lower bound of the concurrence fill, first we choose $\phi=2n\pi$ such that $e^{i\phi}=1$ where $n=1,2,3,...$. Now our aim is to find the value of $p$ for which the state $|\psi_{p,\phi}\rangle$ attains minima. From (37), we have
@@
$p_{1}$ $=$ $\frac{2c^{2}d^{2}-f^{2}b^{2}}{a^{2}b^{2}+2(c^{2}d^{2}-f^{2}b^{2})}$ (45)
$p_{2}$ $=$ $\frac{2c^{2}f^{2}-b^{2}d^{2}}{a^{2}b^{2}+2(c^{2}f^{2}-b^{2}d^{2})}$ (46)
$p_{3}$ $=$ $\frac{2d^{2}f^{2}-b^{2}c^{2}}{a^{2}b^{2}+2(d^{2}f^{2}-b^{2}c^{2})}$ (47)
Now, we find that (44) holds when $|a|^{2}+|b|^{2}=1$ and $c=d=f=\frac{1}{\sqrt{3}}$. Hence, $p_{min}=\frac{1}{1+3b^{2}}$ is the point of minima with $10.25<p_{min}<$ 1, at which the concurrence fill is given by
$C_{F}^{(min)}(|\psi_{p_{min},\phi}\rangle)=\frac{4b^{2}}{1+3b^{2}}\;\text{where}\;0<b<1$ (48)
@@
$C_{F}(|\psi_{p,\phi}\rangle)\geq\frac{4b^{2}}{1+3b^{2}}\;\text{where}\;0<b<1$ (49)
@@
Consider the symmetric GHZ state and symmetric W state in (35). In this case, we have $a=b=\frac{1}{\sqrt{2}}$ and $c=d=f=\frac{1}{\sqrt{3}}$ which gives $p_{min}=\frac{2}{5}$, and the concurrence fill is given as $C_{F}(|\psi_{\frac{2}{5},\phi}\rangle)=\frac{4}{5}$. Hence, the lower bound obtained in (49) conincides with the actual expression of concurrence fill for symmetric mixture of GHZ and W class of states.
@@ IV.2 Concurrence fill for mixed states
The convex roof construction is a widely used method for quantifying the entanglement of mixed states. The ensemble that minimizes the average value of an entanglement measure under this construction is referred to as the optimal ensemble. While the concurrence of two-qubit mixed states admits an analytical solution, in general, determining the convex roof is highly nontrivial. Nonetheless, several effective approaches have been developed to estimate it, including techniques that provide upper and lower bounds as useful approximations oster2008 ; guhne2015 .
Let us consider the mixed state
$\rho=\sum_{i}p_{i}\Omega_{i},\;\Omega_{i}=\frac{|\phi_{i}\rangle\langle\phi_{i}|}{\langle\phi_{i}|\phi_{i}\rangle}$ (50)
@@
$C_{F}(\rho)=min_{\{p_{i},\Omega_{i}\}}\sum_{i}p_{i}C_{F}(\Omega_{i})$ (51)
Our aim is to estimate the concurrence fill of the generalized mixture of GHZ and W class of states. Let us consider the mixed state of the form $\rho(p)$ given in (32). In njp2008 , the optimal decomposition of the zero tangle mixed states $\rho(p)$ has been derived. It is given as follows.
[…]
@@
We have $\tau(\rho(p))=0$ when $0\leq p\leq p_{0}$. The concurrence fill of the state $\hat{\rho}(p)$ is given as
$C_{F}(\hat{\rho}(p))\leq\frac{1}{3}\left[C_{F}(|\psi_{p,0}\rangle)+C_{F}(|\psi_{p,\frac{2\pi}{3}}\rangle)+C_{F}(|\psi_{p,\frac{4\pi}{3}}\rangle)\right]$ (55)
@@
[…]
@@ V Conclusion
In this work, we addressed the challenge of detecting and classifying three qubit genuine entangled states, which remains elusive under many conventional measures despite experimental realizations. To achieve the task, we study the concurrence fill for pure as well as mixed three-qubit states. We began by reformulating the concurrence fill in terms of three tangle and partial tangle. This characterization is important because partial tangle can be measured using expectation value of measurement operators which makes them operationally meaningful. We further employ our formulation to classify GHZ and W class of states. Since convex roof extension is a computationally challenging task, we applied our framework to analyze a rank-2 mixture of generalized GHZ and W-class states, which serves as a representative for nontrivial quantum mixtures. For the eigenstates of the resulting generalized mixed states, we derived an explicit expression for concurrence fill and established an upper bound for the zero tangle mixed states.
These findings not only enhance the utility of concurrence fill as a diagnostic tool for detecting genuine three qubit entangled states, but also open the door to future studies on geometric characterizations of quantum measures in multipartite systems.

### v2

@@ IV Concurrence Fill for a Rank-two Mixture of states
In this section, we study the rank-2 mixture of generalized GHZ and W states. Firstly we derive the minimum value of the concurrence fill for any pure state superposition of generalized GHZ and W states. Next, we formulate the concurrence fill for the mixed states with zero tangle using optimal decomposition method oster2006 ; oster2008 .
@@ IV.1 Concurrence fill for eigenstates of $\rho(p)$
In order to find the concurrence fill for mixed states, let us first study the pure states that are superpositions of the eigenstates of $\rho$. All the elements of the optimal decompositions of $\rho(p)$ are linear combinations of $|gGHZ\rangle$ and $|gW\rangle$. The eigenstates of $\rho(p)$ are given as
$|\psi_{p,\phi}\rangle=\sqrt{p}|gGHZ\rangle-\sqrt{1-p}\;e^{i\phi}|gW\rangle.$ (33)
In oster2006 , authors derived the three-tangle for the pure state $|\psi_{p,\phi}\rangle$ which is given as follows:
@@
Using the deduction obatined in Appendix, we derive the concurrence fill of the state $|\psi_{p,\phi}\rangle$ and it is given as follows:
@@
$C_{F}(|gGHZ\rangle)$ $=$ $4|a^{2}b^{2}|,$ (37)
$C_{F}(|gW\rangle)$ $=$ $8|cdf|\left(\frac{|c^{2}d^{2}+d^{2}f^{2}+c^{2}f^{2}|}{3}\right)^{1/4}.$ (38)
The concurrence fill $C_{F}(|\psi_{p,\phi}\rangle)$ is plotted in Fig-4 for $0\leq p\leq 1$ and different values of $\phi$.
Figure 4: The solid curves represent the concurrence fill of the state $|\psi_{p,\phi}\rangle$ given in (33) for $a=b=\frac{1}{\sqrt{2}}$; $c=d=f=\frac{1}{\sqrt{3}}$ and various values of $\phi$ with $0\leq p\leq 1$.
Interestingly, while the expression for concurrence fill obtained in (35) is simpler to use, it depends on the full set of state parameters. To address this, we derive a lower bound for the concurrence fill specifically for the eigenstates of the rank-two mixture $\rho(p)$. This lower bound is particularly useful, as it depends on only a single state parameter, thereby reducing the number of measurements required to estimate the entanglement in the eigenstates of the mixed state.
Now let us establish the lower bound of the concurrence fill for the pure eigenstates of the form (33), which is appended in the following theorm.
Theorem 3. For the three-qubit pure state $|\psi_{p,\phi}\rangle$ given in (33), the following inequality holds.
$C_{F}(|\psi_{p,\phi}\rangle)\geq C_{F}^{(min)}(|\psi_{p^{*},\phi}\rangle)$ (39)
where
@@
and $C_{F}^{(min)}(|\psi_{p^{*},\phi}\rangle)$ represents the lower bound of the concurrence fill for the state $|\psi_{p^{*},\phi}\rangle$.
@@
Our aim is to find the value of $p$ for which the concurrence fill for the state $|\psi_{p,\phi}\rangle$ attains minima. From (35), we have
@@
$p_{1}$ $=$ $\frac{2e^{4i\phi}c^{2}d^{2}-e^{2i\phi}f^{2}b^{2}}{a^{2}b^{2}+2(e^{4i\phi}c^{2}d^{2}-e^{2i\phi}f^{2}b^{2})}$ (43)
$p_{2}$ $=$ $\frac{2e^{4i\phi}c^{2}f^{2}-e^{2i\phi}b^{2}d^{2}}{a^{2}b^{2}+2(e^{4i\phi}c^{2}f^{2}-e^{2i\phi}b^{2}d^{2})}$ (44)
$p_{3}$ $=$ $\frac{2e^{4i\phi}d^{2}f^{2}-e^{2i\phi}b^{2}c^{2}}{a^{2}b^{2}+2(e^{4i\phi}d^{2}f^{2}-e^{2i\phi}b^{2}c^{2})}$ (45)
Since $p\in[0,1]$ is real, we assume the relative phase $\phi=\frac{n\pi}{2}$ where $n=0,1,2,.$.... Here, the $\pi/2$-periodic behavior stems from the relative phase imparted by the local operation $diag\{\exp(i\pi/2),1\}$ applied individually to every qubit.
Now, we find that the condition (42) holds when $|a|^{2}+|b|^{2}=1$ and $c=d=f=\frac{1}{\sqrt{3}}$. The value of $p^{*}$ satisfying (42) is obtained as (40). Since, $\frac{d^{2}C_{F}(|\psi_{p,\phi}\rangle)}{dp}\geq 0$ for $|a|^{2}+|b|^{2}=1$ and $c=d=f=\frac{1}{\sqrt{3}}$, we conclude that $p^{*}$ obtained in (40) is the point of minima with $0.25<p^{*}<1$.
@@
$C_{F}(|\psi_{p,\phi}\rangle)\geq C_{F}^{(min)}(|\psi_{p^{*},\phi}\rangle)$ (46)
where
[…]
with $0<b<1$. ∎
@@
Consider the symmetric GHZ state and symmetric W state in (33) with $a=b=\frac{1}{\sqrt{2}}$ and $c=d=f=\frac{1}{\sqrt{3}}$. Using (40), we obtain $p^{*}=\frac{2}{5}$. Using (47), the concurrence fill is given as $C_{F}^{(min)}(|\psi_{\frac{2}{5},\phi}\rangle)=\frac{4}{5}$. Hence, the lower bound obtained in (46) conincides with the actual expression of concurrence fill for symmetric mixture of GHZ and W class of states given in (35).
@@ IV.2 Concurrence fill for mixed states
The convex roof construction is a widely used method for quantifying the entanglement of mixed states. Let us consider the mixed state
$\rho=\sum_{i}p_{i}\Omega_{i},\;\Omega_{i}=\frac{|\phi_{i}\rangle\langle\phi_{i}|}{\langle\phi_{i}|\phi_{i}\rangle}.$ (48)
@@
$C_{F}(\rho)=min_{\{p_{i},\Omega_{i}\}}\sum_{i}p_{i}C_{F}(\Omega_{i}).$ (49)
The ensemble that minimizes the average value of an entanglement measure under this construction is referred to as the optimal ensemble. While the concurrence of two-qubit mixed states admits an analytical solution, in general, determining the convex roof is highly nontrivial. Nonetheless, several effective approaches have been developed to estimate it, including techniques that provide upper and lower bounds as useful approximations oster2008 ; guhne2015 .
In njp2008 , authors provide explicit expression of three-tangle for the generalized mixture of GHZ and W class of states $\rho(p)$ given in (30). It is shown that upto a certain value of $p\in[0,1]$, the mixed three-tangle vanishes. Hence, there are large number of states in this mixed state family which have vanishing tangle and hence cannot be quantified using three-tangle. Our aim is to quantify the amount of entanglement for the mixed states $\rho(p)$ with zero tangle. We achieve this by estimating the concurrence fill for these states.
In njp2008 , authors identified the zero polytope containing all the mixed states $\rho(p)$ with zero tangle. The optimal decomposition of these states consists of pure states of the form (33) for which three-tangle is zero. It is shown that the $\tau_{\psi}=0$, for the generalized W state $|gW\rangle$, and three states $|\psi_{p_{0},\frac{2k\pi}{3}}\rangle$ for $n=0,1,2$, where
$p_{0}=\frac{(16c^{2}d^{2}f^{2})^{1/3}}{(a^{4}b^{2})^{1/3}+(16c^{2}d^{2}f^{2})^{1/3}}.$ (50)
We have
$\tau(\rho(p))=0\;\;\text{for}\;\;p\in[0,p_{0}].$ (51)
The optimal decomposition of the zero tangle mixed states $\rho(p)$ is given as follows njp2008 .
$\rho(p)=\frac{p}{p_{0}}\hat{\rho}(p_{0})+\frac{p_{o}-p}{p_{0}}\Omega_{gW};$  $\;0\leq p\leq p_{0}$ (52)
@@
Using (53), the concurrence fill of the state $\hat{\rho}(p)$ is given as
$C_{F}(\hat{\rho}(p))\leq\frac{1}{3}\left[C_{F}(|\psi_{p,0}\rangle)+C_{F}(|\psi_{p,\frac{2\pi}{3}}\rangle)+C_{F}(|\psi_{p,\frac{4\pi}{3}}\rangle)\right].$ (54)
@@
Consider the symmetric GHZ and the symmetric W state, i.e., we have $a=b=\frac{1}{\sqrt{2}}$ and $c=d=f=\frac{1}{\sqrt{3}}$ in (30). Using (50), we obtain $p_{0}=\frac{4\sqrt[3]{2}}{3+4\sqrt[3]{2}}$. Using (56), the upper bound of the concurrence fill for generalized mixture of states represented by $\rho(p)$ for $p\in[0,p_{0}]$ is given in Fig-5.
@@ V Conclusion
In this work, we addressed the challenge of detecting and classifying three qubit genuine entangled states, which remains elusive under many conventional measures despite experimental realizations. To achieve the task, we study the concurrence fill for pure as well as mixed three-qubit states. We began by reformulating the concurrence fill in terms of three tangle and partial tangle. This characterization is important because tangle is LOCC monotonic and partial tangle can be measured using expectation value of measurement operators which makes them operationally meaningful. We further employ our formulation to classify GHZ and W class of states. Since convex roof extension is a computationally challenging task, we applied our framework to analyze a rank-2 mixture of generalized GHZ and W-class states, which serves as a representative for nontrivial mixed quantum states. For the eigenstates of the resulting generalized mixed states, we derived an explicit expression for concurrence fill and established an upper bound for the zero tangle mixed states. These findings not only enhance the utility of concurrence fill as a diagnostic tool for detecting genuine three qubit entangled states, but also open the door to future studies on geometric characterizations of quantum measures in multipartite systems.
@@ VII DATA AVAILABILITY STATEMENT
VIII Appendix
Consider the pure three-qubit state $|\psi_{ABC}\rangle$. Let $\rho_{AB}$ and $\rho_{AC}$ denote the density matrices representing the reduced states of the three-qubit state $\rho_{ABC}=|\psi_{p,\phi}\rangle\langle\psi_{p,\phi}|$. Let $\tilde{\rho}_{ij}$ be the spin flipped density matrix as defined in (4) and $\lambda_{n}^{ij}$ be the square root of the $n^{th}$ eigenvalue of the state $\rho_{ij}{\tilde{\rho}}_{ij}$, where $i,j$ are the distinct index representing the subsystems $A,B,C$ and eigenvalues are arranged in decreasing order. In ckw2000 , authors defined the concurrence of the reduced two-qubit state as follows:
$C_{ij}=(\lambda_{1}^{ij}-\lambda_{2}^{ij})\;\text{for}\;i\neq j.$ (57)
The three-tangle defined in (8) can be defined in terms of the eigenvalues of reduced states as follows ckw2000 .
$\tau(|\psi_{p,\phi}\rangle)=2(\lambda_{1}^{AB}\lambda_{2}^{AB}+\lambda_{1}^{AC}\lambda_{2}^{AC})$ (58)
Using (57) and (58) in (8), we obatin
$C^{2}_{A|BC}$ $=$ $\tau(|\psi_{p,\phi}\rangle)+C^{2}_{AB}+C^{2}_{AC}$ (59)
$=$ $2(\lambda_{1}^{AB}\lambda_{2}^{AB}+\lambda_{1}^{AC}\lambda_{2}^{AC})$
$+(\lambda_{1}^{AB}-\lambda_{2}^{AB})^{2}+(\lambda_{1}^{AC}-\lambda_{2}^{AC})^{2}$
$=$ ${\lambda_{1}^{2}}^{(AB)}+{\lambda_{2}^{2}}^{(AB)}+{\lambda_{1}^{2}}^{(AC)}+{\lambda_{2}^{2}}^{(AC)}$
Similarly, we have
$C^{2}_{B|AC}$ $=$ ${\lambda_{1}^{2}}^{(AB)}+{\lambda_{2}^{2}}^{(AB)}+{\lambda_{1}^{2}}^{(BC)}+{\lambda_{2}^{2}}^{(BC)}$ (60)
and
$C^{2}_{C|AB}$ $=$ ${\lambda_{1}^{2}}^{(BC)}+{\lambda_{2}^{2}}^{(BC)}+{\lambda_{1}^{2}}^{(AC)}+{\lambda_{2}^{2}}^{(AC)}$ (61)
Using above equations (59)-(61) in (12), we obtain
$Q=\sum_{k=1,2}({\lambda_{k}^{2}}^{(AB)}+{\lambda_{k}^{2}}^{(BC)}+{\lambda_{k}^{2}}^{(AC)})$ (62)
and hence the concurrence fill is obtained as
$C_{F}(|\psi_{ABC}\rangle)=\bigg{(}\frac{16}{3}\sum_{k=1,2}({\lambda_{k}^{2}}^{(AB)}+{\lambda_{k}^{2}}^{(BC)}+{\lambda_{k}^{2}}^{(AC)})\\ \times({\lambda_{1}^{2}}^{(AB)}+{\lambda_{2}^{2}}^{(AB)})({\lambda_{1}^{2}}^{(BC)}+{\lambda_{2}^{2}}^{(BC)})({\lambda_{1}^{2}}^{(AC)}+{\lambda_{2}^{2}}^{(AC)})\bigg{)}^{1/4}.$ (63)